Consider a random variable X with the following probability mass function P(X=0)=0.25, P(X=5)=0.5, P(X=12)=0.25. What is the expected value (or mean) of X?
Answer #1

Here P(X = 0) P(X-5) P(X-12) = 0.25 + 0.5 0.25 1

So expected value (or mean) of X= Er.P(X 2) = 0(0.25) 5(0.5) 12(0.25) 5.5